find area of a rectangle if its perimeter is 30 and its base is 8
Answer by RAY100(1637)

Perimeter = 2* Length +2* Width
But L = 8,,,,and P=30
P=2L + 2W = 30 = 2*8 + 2W,,,,W = 7
Area = L * W = 8 * 7 = 56